The area of a rectangular wall of a barn is 96 square feet. Its length is 4 feet longer than twice its width. Find the length and width of the wall of the barn.

Answer by Boreal(15235)

A=l*w
l=2w+4
substitute
A=w*(2w+4)=2w^2+4w=96
divide by 2 and subtract the constant from both sides
w^2+2w-48=0
(w+8)(w-6)=0
only positive root is width=6 feet and length=16 feet.